John invested $7000 into two funds. One fund paid 14% interest per year and the other paid 11% interest per year. The total interest for the year was $902. How much money was invested at each rate?
Answer by jayanthihemkos@yahoo.com(87)

John invested $7000 into two funds. One fund paid 14% interest per year and the other paid 11% interest per year. The total interest for the year was $902. How much money was invested at each rate?
Let one fund = x
The other fund = 7000-x
Interest paid for first fund = 14% of x = (14/100)x
Interest paid for the second fund = 11% of (7000-x) = (11/100)(7000-x)
Total interest = $902
(14/100)x +(11/100)(7000-x) = $902
(14x + 77000-11x)/100 = $902

(3x+77000)/100 = $902

(3x+77000) /100*100 = 902*100
3x +77000 = 90200
3x +77000-77000 = 90200-77000
3x = 13200
3x/3 = 13200/3
x = $4400--- interest paid for the first fund
7000-x = 7000-4400 = $2600 Interest paid for the second fund
First fund interest = $4400
Second fund interest = $2600